# Epidemiology of Large Language Models: A Benchmark for Observational Distribution Knowledge

### v1

@@ Appendix F Does Fine-Tuning Improve Performance?
In this appendix, we evaluate whether fine-tuning helps the performance of LLMs on our benchmark. Given the relatively poor performance of models across a range of tasks, a natural question arises whether additional training could improve models’ knowledge of observational distributions. We investigate this hypothesis by fine-tuning a LLaMA3 8B instruction-tuned model on synthetic data generated from our ground truth datasets, and comparing the performance to the baseline pretrained model.
F.1 Synthetic Data Creation
@@
Our first step is to create a synthetic dataset that simulates the type of text a model might be given access to during its pretraining stage. To generate such that, we use a pretrained Gemma3 27B instruction model (team2025gemma) and focus on the NSDUH data. Specifically, we sample different individuals from the dataset, reflecting the observational distribution. Upon selecting an individual, we generate a text paragraph representing the person and their associated covariates. For NSDUH data, different tasks look at cigarette, alcohol, marijuana, cocaine, and heroin usage across demographics and therefore each generated textual note must contain all of this information. In particular, we use the following prompt template for querying the Gemma3 27B model:
@@
You are a data generator. Follow the rules strictly.
RULES:
@@
<story>
(your narrative here)
</story>
We create a dataset of 2,500 samples. An example of a generated data sample is shown below, showcasing the type of synthetic data to be used for fine-tuning:
@@
F.2 Fine-Tuning
By fine-tuning a decoder-only model (LLama3 8B or Mistral 7B) on the above dataset, we test whether a model can internalize observational knowledge by learning to do causal language modeling on narrative text. An alternative fine-tuning approach would be to update model weights directly on summarized statistics. However, this way of training would no longer represent embedding probabilistic knowledge (which our benchmark aims to assess) into the model, and would instead embed factual knowledge.
@@ F.2 Fine-Tuning
As the loss behavior indicates, the fine-tuning approach reduces both training and evaluation losses early on, after which overfitting begins, and the training loss continues to decrease while evaluation loss starts to increase. Model checkpoint saved using early stopping is indicated with a vertical red line.
F.3 Evaluation & Observations
@@
In the last step, we evaluate the fine-tuned models on all of the NSDUH tasks, and compare the task performances to those of pretrained LLama3 8B and Mistral 7B models. The results are shown in Fig. LABEL:fig:ft-results for both settings (low- and high-dimensional) and prompting techniques (question-answer and likelihood prompting).
@@
Notably, for Q&A prompting, fine-tuning does not seem to improve performance across either low- or high-dimensional settings – very few and modest score increases are recorded for both models. At the same time, in the high-dimensional setting, for some tasks where a pretrained LLama3 8B model scored above 0 the performance drops with fine-tuning.
@@
For likelihood prompting, the impact of fine-tuning is different for low-dimensional and high-dimensional settings. In the low-dimensional setting, FT increases the performance on 6 out of 10 tasks for LLama3 8B and 3 out of 10 for Mistral 7B. In the high-dimensional setting, the performance is not improved on any of the 33 tasks for either model, so both the pretrained and fine-tuned models score exactly 0 on each task.
@@
Therefore, some FT gains are observed for likelihood prompting in the low-dimensional setting. Interestingly, one would perhaps hypothesize that fine-tuning on the task of causal language modeling would have a greater impact and improve upon performance for the Q&A prompting, since this prompting technique more directly elicits the model’s internal probability from next token predictions (instead of asking for a likelihood). However, the results do not show this. Therefore, our investigation in this appendix offers preliminary evidence showing that fine-tuning models on text generated from the correct observational distribution does not ensure that models internalize knowledge of these distributions. Future work is needed to extend these findings to other models and data sources, and to investigate methods which may improve the abilities of models in terms of probabilistic L1 knowledge.

### v2

@@ Appendix C Does Fine-Tuning Improve Performance?
In this appendix, we evaluate whether fine-tuning helps the performance of LLMs on our benchmark. Given the relatively poor performance of models across a range of tasks, a natural question arises whether additional training could improve models’ knowledge of observational distributions. We investigate this hypothesis by fine-tuning a LLaMA3 8B and Mistral 7B instruction-tuned models on synthetic data generated from our ground truth datasets, and comparing the performance to the baseline pretrained models.
C.1 Synthetic Data Creation
@@
Our first step is to create a synthetic dataset that simulates the type of text a model might be given access to during its pretraining stage. To generate such that, we use a pretrained Gemma3 27B instruction model (team2025gemma) and focus on the NSDUH and BRFSS data. Specifically, we sample different individuals from the dataset, reflecting the observational distribution. Upon selecting an individual, we generate a text paragraph representing the person and their associated covariates.
C.1.1 NSDUH Data
For NSDUH data, we use the following prompt template for querying the Gemma3 27B model:
@@ C.1.1 NSDUH Data
You are a data generator. Follow the RULES strictly:
@@
<story> (your narrative here) </story>
Such a prompt is used since tasks look at cigarette, alcohol, marijuana, cocaine, and heroin usage across demographics and therefore each generated textual note must contain all of this information. In particular, We create a dataset of 2,500 samples. An example of a generated data sample is shown below, showcasing the type of data to be used for fine-tuning:
@@
C.1.2 BRFSS Data
For the BRFSS data, we use a very similar prompt template for querying the Gemma3 27B model, but change the facts generated by the model:
Prompt
FACTS:
- state: {state}
- age: {age}
- sex: {sex}
- race: {race}
- education: {education}
- income: {income}
- smoking status: {smoker}
- BMI: {bmi}
- exercise monthly: {exercise_monthly}
- poor mental health: {poor_mental_health}
- diabetes: {diabetes}
- high blood pressure: {high_bp}
- asthma: {asthma}
- cholesterol: {cholesterol}
- heart attack: {heart_attack}
- stroke: {stroke}
- depression: {depression}
Since different tasks look at health status across demographics, so the textual data must include all of these attributes. Again, we create a dataset of 2,500 samples, such as:
Example Output Eleanor, a 52.0 year old White female, lived a relatively quiet life in Ohio. She was a college graduate earning between \$50–100k annually. She had never been diagnosed with diabetes, high blood pressure, asthma, or high cholesterol. Thankfully, she’d also avoided both a heart attack and stroke. Eleanor did not smoke and maintained a BMI of 25.84, though she admitted her exercise routine was limited to about monthly activity. She reported good physical health and fortunately, had no history of depression. Her mental health was stable; she didn’t suffer from poor mental health.
C.2 Fine-Tuning
By fine-tuning a decoder-only model (LLama3 8B or Mistral 7B) on the above two datasets, we test whether a model can internalize observational knowledge by learning to do causal language modeling on narrative text. An alternative fine-tuning approach would be to update model weights directly on summarized statistics. However, this way of training would no longer represent embedding probabilistic knowledge (which our benchmark aims to assess) into the model, and would instead embed factual knowledge.
@@ C.2 Fine-Tuning
As the loss behavior indicates, the fine-tuning approach reduces both training and evaluation losses early on, after which overfitting begins, and the training loss continues to decrease while evaluation loss starts to increase. This observation is consistent across models and datasets, showing adequate fine-tuning behavior. Model checkpoints saved using early stopping are indicated with a vertical red line.
C.3 Evaluation & Observations
@@
In the last step, we evaluate the fine-tuned models on all of the NSDUH and BRFSS tasks, and compare the task performances to those of pretrained LLama3 8B and Mistral 7B models. The results are shown in Fig. LABEL:fig:ft-results and Fig. LABEL:fig:ft-results-brfss for both settings (low- and high-dimensional) and prompting techniques (question-answer and likelihood prompting) for the respective datasets.
C.3.1 NSDUH Data
@@
For QA prompting, fine-tuning does not seem to improve performance across either low- or high-dimensional settings – very few and modest score increases are recorded for both models. At the same time, in the high-dimensional setting, for some tasks where a pretrained LLama3 8B model scored above 0 the performance drops with fine-tuning.
@@
For likelihood prompting, the impact of fine-tuning is different for low-dimensional and high-dimensional settings. In the low-dimensional setting, FT increases the performance on 6 out of 10 tasks for LLama3 8B and 3 out of 10 for Mistral 7B. In the high-dimensional setting, the performance is not improved on any of the 33 tasks for either model, so both the pretrained and fine-tuned models score exactly 0 on each task. Therefore, some FT gains are observed for likelihood prompting in the low-dimensional setting.
C.3.2 BRFSS Data
For QA prompting, fine-tuning leads to some performance improvements (particularly for Mistral 7B, see Fig. LABEL:fig:ft-mistral-brfss), but these do not appear consistent across tasks. Furthermore, there are also instances of a drop in performance in certain tasks. For Llama3 8B the results are also not conclusive, with some improvements on tasks where the pretrained model score 0, but also a drastic drop in performance across tasks where the pretrained model scored more favorably.
For likelihood prompting, the impact of fine-tuning is again different for low-dimensional and high-dimensional settings. In the low-dimensional setting, FT increases the performance on 4 out of 10 tasks for LLama3 8B, yet it also reduces the performance to zero on one task. Similarly, for Mistral 7B FT reduces the performance on 3 out of 10 tasks, and show an increase for one task. In the high-dimensional setting, the performance is not improved on any of the 33 tasks for either model; in fact the fine-tuned Mistral 7B model performs worse than the pretrained one, and has a consistent 0 score across all tasks. Thus, no consistent improvements are observed across different settings.
C.3.3 Fine-Tuning Takeaways
@@
One would perhaps hypothesize that fine-tuning on the task of causal language modeling would have a greater impact and improve upon performance for the QA prompting, since this prompting technique more directly elicits the model’s internal probability from next token predictions (instead of asking for a likelihood), and the fine-tuning data encodes the correct underlying distribution in this form. However, the observed results do not support this hypothesis. Therefore, our investigation in this appendix offers preliminary evidence that fine-tuning models on text generated from the correct observational distribution does not ensure that models internalize knowledge of these distributions. Future work is needed to extend these findings to other models and data sources, and to investigate methods that may improve the abilities of models in terms of probabilistic L1 knowledge.
Appendix D Fairness Impacts
In this appendix, we assess whether the presence of protected attributes in a task influences model performance. Language models are often trained with a variety of safety-oriented or debiasing interventions, e.g., safety tuning, data mixture filtering, or RLHF reward shaping (bolukbasi2016man; zhao2018learning; glaese2022improving). These techniques may cause models to behave differently when a prompt mentions sensitive attributes such as sex or race. If such mechanisms are active, they could in principle deflate scores on tasks where protected attributes appear directly in the question; hence it is important to verify that the main results are not driven by this effect. To examine this, we partition the benchmark tasks into two groups: (i) tasks that reference attributes sex or race (118/169 tasks), and (ii) tasks that contain no such attributes (51/169 tasks). For each model and prompting mode, we compute average performance separately in these two groups.
The results are shown in Fig. LABEL:fig:fairness-impact, with a split in sensitive and non-sensitive tasks. Overall, most models perform slightly better on non-sensitive tasks, although not all. Furthermore, the performance difference appears to be modest, and the overall pattern of findings remains unchanged. These results support the fact that the benchmark’s findings are not driven purely by the debiasing techniques applied on protected attributes. Tasks involving sex or race exhibit slightly worse scores, but the qualitative picture of model underperformance persists across both categories.
Appendix E Data Contamination
In this appendix, we analyze potential data contamination during pretraining, that is, whether during pretraining models may have seen the exact questions used in the benchmark. There is a non-zero chance for this, as language models are trained on massive text corpora. As discussed in (hendrycks2020measuring), if a model simply recalls a memorized question-answer pair, this would be reflected in lower question entropy and a higher score. Such data contamination would also make reported model performance appear more favorable than the true underlying model capability.
In light of this, we investigate whether there is a relationship between question entropy and task performance. We focus on likelihood prompting, as this is the more likely question format that models may have encountered during training. In Fig. LABEL:fig:data-contamination, we plot the score and question entropy for each low-dimensional and high-dimensional task for LLaMA3 8B and Phi-4 models. As the figure illustrates, there is an absence of a positive correlation between score and question entropy, suggesting that there is no evidence that exact question–answer pairs were memorized during pretraining. For a further discussion on investigating question memorization using entropy, we refer the reader to (hendrycks2020measuring) and (brown2020language).